OpenStudy (anonymous):

Solve for x: (x – 4)2 = 2

OpenStudy (anonymous):

2x-8=2 2x=10 x=5

OpenStudy (anonymous):

is it multiplied by 2?

OpenStudy (hba):

do u know (a-b)^2=a^2-2ab+b^2 i hope u can solve it now

OpenStudy (anonymous):

if it is raised by 2, then you can simply solved x-4=2 x=6 multiple 2

OpenStudy (jiteshmeghwal9):

use identity\[(a-b)^2=a^2+b^2-2ab\]so,\[(x^2-8x+16)=2\]\[x^2-8x+16-2=0\]\[x^2-8x+14=0\]

OpenStudy (jiteshmeghwal9):

now use quadratic formula\[-b \pm \sqrt{b^2-4ac}\over2a\]

OpenStudy (jiteshmeghwal9):

\[(-8) \pm \sqrt{(-8)^2-4(1)(14)}\over 2(1)\]\[8 \pm \sqrt{64-56}\over2\]\[8 \pm \sqrt{8}\over 2 \]\[\alpha={8+\sqrt{8}\over2}\]\[\beta={8-\sqrt{8}\over2}\]

OpenStudy (anonymous):

( x - 4 ) 2 = 2 divide both sides by 2 x - 4 = 0 x = 4 thats the best answer i could give out

OpenStudy (anonymous):

x = 4